A car speedometer has a 4% uncertainty. What is the range of possible speeds (in km/h) when it reads 60 km/h?
Answer #1

Solution:

Uncertainty = 4 * 60 / 100 = 2.4 km/h

=> range = 60 - 2.4 to 60+2.4

=> rand = 57.6 - 62.4 km/h

please comment for queries.

please rate thanks.